# Distribution of lengths of closed saddle connections on moduli space of large genus translation surface

### v1

@@ 4.2.3. Pinching
$\operatorname{Vol}(\mathcal{C}^{1,0}_{g}(2^{g-1})))=\operatorname{Vol}(\mathcal{C}^{1,0}_{g,Per}(2^{g-1})))=\frac{\pi(b^{2}-a^{2})}{g}\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2})).$
5. Collapse closed saddle connections simultaneously
From the construction above we can collapse one closed saddle connection on some $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$ and obtain a new surface $(X^{\prime},\omega^{\prime})\in{\mathcal{H}}_{g-1}(2^{g-2},0,0)$. If we want to collapse any $k$ closed saddle connections, we have to ensure there aren’t two short closed saddle connections sharing one zero. We need the following collar lemma [12, 6, 9].
5.1. Collar lemma
This section we will introduce the collar lemma on flat surface, and use it to prove the closed saddle connections we want to collapse are disjoint.
Definition 5.1.
Let $A$ be an annulus on $(X,\omega)$ such that there are no zeros of $\omega$ in the interior. We say that $A$ is expanding if the curvature of each boundary component of $A$ has constant sign, either positive or negative at each point, and the boundary curves are equidistant.
Denote the boundary components by $\partial_{0}A$ and $\partial_{1}A$. If $A$ is an expanding annulus, then we assume that the total curvature $\kappa(\partial A_{0})=\kappa(A)>0$ and $\kappa(\partial A_{1})=-\kappa(A)<0$. Let $d(A)$ be the distance between $\partial_{0}A$ and $\partial_{1}A$.
For $\gamma$ a simple closed curve , let $C(\gamma)$ be the maximal flat cylinder on $(X,\omega)$ with core curves homotopic to $\gamma$, and let $A(\gamma)$ be the maximal expanding annulus, that is, one of whose boundary components is a $\omega$-geodesic homotopic to $\gamma$.
Lemma 5.2.
[9, Lemma 1] Assume that $\gamma$ is a simple closed curve on $(X,\omega)$ such that its hyperbolic length $\ell_{X}(\gamma)$ is sufficiently small. Then
  1. (1)
    $\frac{1}{\ell_{X}(\gamma)}\asymp\max\{\operatorname{Mod}(A(\gamma)),\operatorname{Mod}(C(\gamma))\};$
  2. (2)
    $\operatorname{Mod}(A(\gamma))\asymp\log\frac{d(A(\gamma))}{\ell_{\omega}(\gamma)}.$
This lemma originates from Minsky [12, §4.3], moreover from the proof of Minsky we know in fact the comparability coefficient depends on the total curvature $\kappa(A(\gamma))$.
For simplicity, denote by $A=A(\gamma)$ and $C=C(\gamma)$. Either $\frac{1}{\ell_{X}(\gamma)}\asymp\operatorname{Mod}(C)$ or $\frac{1}{\ell_{X}(\gamma)}\asymp\operatorname{Mod}(A)$. In the first case, we have $\operatorname{Mod}(C)=\frac{d(C)}{\ell_{\omega}(\gamma)}$, where $d(C)$ denotes the width of $C$. Thus we have
(5.1) $d(C)\asymp\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}.$
In the second case, we have
(5.2) $\log\frac{d(A)}{\ell_{\omega}(\gamma)}\asymp\operatorname{Mod}(A)\asymp\frac{1}{\ell_{X}(\gamma)},$
which implies
$d(A)\succ\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}.$
We conclude that there is an annulus homotopic to $\gamma$ such that if $\ell_{X}(\gamma)$ is sufficiently small then the width of the annulus is at least
$c\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)},$
where $c>0$ is a constant depending on $\kappa(A(\gamma))$.
@@ 5.1. Collar lemma
For $g$ large enough, let $\gamma_{1},\cdots,\gamma_{k}$ be $k$ closed saddle connections of multiplicity $1$ with no cylinder around them on $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$. If their length are all in $\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}]$, $\gamma_{1},\cdots,\gamma_{k}$ are disjoint.
@@
By assumption, any $\gamma_{i}$ has only one geodesic representation. Since the minimal intersection between two homotopic class is realized by its geodesic representation, if $\gamma_{i}$ and $\gamma_{j}$ intersect, they are intersected essentially. From the collar lemma 5.2, $\gamma_{i}$ must go through the collar (in fact, it’s expanding) $A_{j}$ around $\gamma_{j}$. As $\gamma_{i}$ and $\gamma_{j}$ have only one double zero on them, their total curvatures have a uniform bound. Then the width of $A_{j}$ has a lower bound $c\frac{\ell_{\omega}(\gamma)}{\ell_{X}(\gamma)}$, where $c$ is a constant. When $g$ goes to infinity, $\ell_{\omega}(\gamma_{i})$ goes to zero, and $\ell_{X}(\gamma_{i})$ must go to zero. So for $g$ large enough, we have
$c\frac{\ell_{\omega}(\gamma_{i})}{\ell_{X}(\gamma)}\geq c\frac{a}{\sqrt{g}\ell_{X}(\gamma)}\geq\frac{b}{\sqrt{g}},$
and we have $\ell_{\omega}(\gamma_{j})\geq\frac{b}{\sqrt{g}}$, which is a contradiction. ∎
@@ 5.2. Collapsing simultaneously
Define $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1}))=(X,\omega,\Gamma_{1},\cdots,\Gamma_{k})$, where $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1}))$, and $\Gamma_{i}$ is an ordered list of $n_{i}$ closed saddle connections $(\gamma_{1,i},\cdots,\gamma_{n_{i},i})$ with lengths in $\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]$. Suppose $a_{1}\leq b_{1}\leq a_{2}\leq b_{2}\cdots\leq a_{k}\leq b_{k}$. Note that when choose $a=a_{1},b=b_{k}$, by Proposition 5.3, these closed saddle connections are disjoint when $g\to\infty$.
We will adjust the surgery from last section to collapse the closed saddle connections in $\Gamma_{1},\cdots,\Gamma_{k}$ in order. The opening and pinching operations are same to the situation of one curve. The only modification is the equation 4.1: by remark 4.3, we need to choose the moving vector $\upsilon$ to ensure the area unchanged, hence we can choose a basis containing $\gamma^{\prime}_{1},\gamma_{1}\cdots,\gamma^{\prime}_{K},\gamma_{K}$, where $\gamma^{\prime}_{i}$ is a saddle connection from opening $\gamma_{i}$. Then the moving vector $\upsilon$ can be constructed like equation 4.1: after defining the value on $\gamma_{i}$, we can choose another relative homotopy class and give a value of $\upsilon$ on it to ensure the area unchanged. This means we open and move the closed saddle connections simultaneously and then pinch them.
@@
$\hat{F}:\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1}))\to M_{K}\cdot{\mathcal{H}}_{g-K}(2^{g-1-K})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}}$
where $M_{K}$ is a combinatorial data consisting of choosing $K$ ordered zeros from $g$ zeros to label. As the discussion above, the mapping is one-to-one and measure-preserving. So we have
$\int_{{\mathcal{H}}_{g}(2^{g-1})}(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}d\mu_{MV}$ $=\operatorname{Vol}(\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})))$
@@
where the limit is because $|M_{K}|=g\cdots(g-K+1)$ and
$\lim_{g\to\infty}\frac{g\cdots(g-K+1)}{g^{n_{1}+\cdots+n_{k}}}=\frac{g\cdots(g-K+1)}{g^{K}}=1.$
Then the limit of expectation is
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1-K}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}$
From Theorem 2.2 ,
@@
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}[3\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
By theorem 2.5, we get theorem 1.2.
From the proof above we can prove Corollary 1.4.
proof of Corollary 1.4.
Since we don’t need to deal with simple zeros in the collapsing surgery, samilarly to the proof of theorem 1.2, we have
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{O(g)-K},1^{2g-2-2O(g)}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{O(g)},1^{2g-2-2O(g)}))}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}},$
and
$\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1},1^{2g-2}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1},1^{2g-2}))}=3^{K}.$
Then
$\lim_{g\to\infty}\mathbb{E}[(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}[3\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$

### v2

@@ 5.1. The measure of exception set
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=0.$
Proof.
For $(X,\omega)\in\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}})$, let $p$ be a zero on $(X,\omega)$ and $\gamma_{1},\gamma_{2}$ be the closed saddle connections with lengths less than $\frac{B}{\sqrt{g}}$ at $p$.
As above we can collapse $\gamma_{1}$ and get a new translation surface in ${\mathcal{H}}_{g-1}(2^{g-2},0,0)$. After the surgery $\gamma_{2}$ will become a segment connecting two marked regular points which are from the pinching surgery. So the image of mapping $F_{3}$ on $\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}})$ is in ${\mathcal{H}}_{g-1,\frac{B}{\sqrt{g}}}(2^{g-2},0,0)\times D(\frac{B}{\sqrt{g}})$, where ${\mathcal{H}}_{g-1,\frac{B}{\sqrt{g}}}(2^{g-2},0,0)$ is the subset that there exists a segment connnecting the two marked points of length less than $\frac{B}{\sqrt{g}}$ and $D(\frac{B}{\sqrt{g}})$ is the disk of radius $\frac{B}{\sqrt{g}}$. From [1, Theorem 1.2], the Siegel-Veech constant of saddle connections connecting two fix zeros of order $m_{1}$ and $m_{2}$ is
$c=(m_{1}+1)(m_{2}+1)(1+O(\frac{1}{g})).$
So we have
$\operatorname{Vol}(\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))\leq(g-1)\frac{B^{2}}{g}\operatorname{Vol}({\mathcal{H}}_{g-1,\frac{B}{\sqrt{g}}}(2^{g-2},0,0))\leq c(g-1)\frac{B^{4}}{g^{2}}\operatorname{Vol}({\mathcal{H}}_{g-1}(2^{g-2},0,0)),$
where the coefficient g1g-1 is the choice of $p$. Then by 2.2, we have
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}\leq\lim_{g\to\infty}3c\frac{B^{4}}{g}=0$
For $\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})$ we also have
@@
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=0.$
@@
For $(X,\omega)\in\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})$, let $p_{1},p_{2}$ be the two zeros on $(X,\omega)$ with distance less than $\frac{B}{\sqrt{g}}$ and $\gamma_{1},\gamma_{2}$ be the two closed saddle connections connecting $p_{1},p_{2}$. For every $k\in\mathbb{N}^{+}$, denote by $\mathcal{C}_{g}^{1,k}(\frac{B}{\sqrt{g}})$ the subset of $\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})$ that the geodesic between $p_{1}$ and $p_{2}$ is a concatenation of $k$ open saddle connections. Their lengths are less than $\frac{B}{\sqrt{g}}$. Choose the shortest saddle connection, using the surgery in [9] we can collapse it and get a new translation surface in ${\mathcal{H}}_{g}(3,1,2^{g-3})$. This is because the locus, that is the extension of the shortest saddle connection can not intersect the whole geodesic. Repeat the surgery until $p_{1}$ and $p_{2}$ are collapsed to one zero and we get a map
$F:\mathcal{C}_{g}^{1,k}(\frac{B}{\sqrt{g}})\to\mathcal{C}_{g}^{2}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}})\times D^{k}(\frac{B}{\sqrt{g}}),$
where $\mathcal{C}_{g}^{2}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}})$ is the subset of ${\mathcal{H}}_{g}(2+k,1^{k},2^{g-2-k})$ such that there are two closed saddle connections $\gamma_{1},\gamma_{2}$ connecting the zero of order $2+k$ with lengths less than $\frac{B}{\sqrt{g}}$ .
Similarly one can collapse $\gamma_{1}$ and the other will become a saddle connection connecting two zeros. Note that the order is larger than $2$, although the collapsing surgery can also be realized, we need to require the angles on the two sides of the closed saddle connections collapsed, which decide the order of zeros after collapsing. We will explain this situation in detail in Section 6.
Let $\mathcal{C}_{g}^{2,b^{\prime},b^{\prime\prime}}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}})$ be the subset of $\mathcal{C}_{g}^{2}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}})$ such that the two angles on both sides of $\gamma_{1}$ are $(2b^{\prime}+1)\pi$ and $(2b^{\prime\prime}+1)\pi$, where $b^{\prime}+b^{\prime\prime}=2+k$. Then after collapsing we get a map
$F^{\prime}:\mathcal{C}_{g}^{2,b^{\prime},b^{\prime\prime}}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}})\to{\mathcal{H}}_{g-1,\frac{B}{\sqrt{g}}}(b^{\prime}-1,b^{\prime\prime}-1,1^{k},2^{g-2-k})\times D(\frac{B}{\sqrt{g}}).$
Similar to Proposition 5.1, we have
$\operatorname{Vol}(\mathcal{C}_{g}^{2,b^{\prime},b^{\prime\prime}}(2+k,1^{k},2^{g-2-k})(\frac{B}{\sqrt{g}}))\leq b^{\prime}b^{\prime\prime}\frac{B^{4}}{g^{2}}\operatorname{Vol}({\mathcal{H}}_{g-1}(b^{\prime}-1,b^{\prime\prime}-1,1^{k},2^{g-2-k})).$
Sum all $k$ and $(b^{\prime},b^{\prime\prime})$ we have
$\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}}))\leq(g-1)(g-2)\sum_{k}\sum_{(b^{\prime},b^{\prime\prime})}b^{\prime}b^{\prime\prime}\frac{B^{4}}{g^{2}}\operatorname{Vol}({\mathcal{H}}_{g-1}(b^{\prime}-1,b^{\prime\prime}-1,1^{k},2^{g-2-k}))(\frac{B^{2}}{g})^{k},$
where the coefficient (g1)(g2)(g-1)(g-2) is the choice of $p_{1}$ and $p_{2}$. Then when $g\to\infty$ we have
$\frac{\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}\leq\sum_{k}\sum_{(b^{\prime},b^{\prime\prime})}b^{\prime}b^{\prime\prime}B^{4}\frac{\operatorname{Vol}({\mathcal{H}}_{g-1}(b^{\prime}-1,b^{\prime\prime}-1,1^{k},2^{g-2-k}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}(\frac{B^{2}}{g})^{k}=O(\frac{1}{g}).$
Combine the two proposition we have
Corollary 5.3.
$\frac{\operatorname{Vol}(\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a}{\sqrt{g}},\frac{b}{\sqrt{g}}])}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=O(\frac{1}{g}).$
@@ 5.2. Collapsing simultaneously
Define $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})=(X,\omega,\Gamma_{1},\cdots,\Gamma_{k})$, where $(X,\omega)\in\mathcal{C}^{1,0}_{g}(2^{g-1})$, and $\Gamma_{i}$ is an ordered list of $n_{i}$ closed saddle connections $(\gamma_{1,i},\cdots,\gamma_{n_{i},i})$ with lengths in $\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]$. Suppose $a_{1}\leq b_{1}\leq a_{2}\leq b_{2}\cdots\leq a_{k}\leq b_{k}$, let $B=b_{k}$. Then we have
$\bigcup_{i}\mathcal{C}^{exc}_{g,Per}(2^{g-1})[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]\subset\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}).$
Let $\mathcal{C}^{\prime}_{g,Per}=\mathcal{C}^{1,0}_{g,Per}(2^{g-1})\setminus(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))$, we will adjust the surgery from last section on $\mathcal{C}^{\prime}_{g,Per}$ to collapse the closed saddle connections in $\Gamma_{1},\cdots,\Gamma_{k}$ in order. The opening and pinching operations are same to the situation of one curve. The only modification is the equation 4.1: by remark 4.3, we need to choose the moving vector $\upsilon$ to ensure the area unchanged, hence we can choose a basis containing $\gamma^{\prime}_{1},\gamma_{1}\cdots,\gamma^{\prime}_{K},\gamma_{K}$, where $\gamma^{\prime}_{i}$ is a saddle connection from opening $\gamma_{i}$. Then the moving vector $\upsilon$ can be constructed like equation 4.1: after defining the value on $\gamma_{i}$, we can choose another relative homotopy class and give a value of $\upsilon$ on it to ensure the area unchanged. This means we open and move the closed saddle connections simultaneously and then pinch them.
@@
$\hat{F}:\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})\to M_{K}\cdot\tilde{{\mathcal{H}}}_{g-K}(2^{g-1-K},0^{2K})\times\prod^{k}_{i=1}(A_{[\frac{a_{i}}{\sqrt{g}},\frac{b_{i}}{\sqrt{g}}]})^{n_{i}},$
where $\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})$ is defined as $\hat{\mathcal{C}}^{1,0}_{g}(2^{g-1})$ and $M_{K}$ is a combinatorial data consisting of choosing $K$ ordered zeros from g1g-1 zeros to label. As the discussion above, we have
$\int_{\hat{\mathcal{C}}^{\prime}_{g,Per}(2^{g-1})}(L_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L_{g,[a_{k},b_{k}]})_{n_{k}}d\mu_{MV}$
$=\operatorname{Vol}(\hat{\mathcal{C}}^{\prime}_{g}(2^{g-1})))$
@@
where the limit is because $|M_{K}|=(g-1)\cdots(g-K)$ and
$\lim_{g\to\infty}\frac{(g-1)\cdots(g-K)}{g^{n_{1}+\cdots+n_{k}}}=\frac{(g-1)\cdots(g-K)}{g^{K}}=1.$
Let $L^{\prime}_{g,[a_{i},b_{i}]}$ be the restriction of $L_{g,[a_{i},b_{i}]}$ on $\mathcal{C}^{\prime}_{g,Per}(2^{g-1})$. Then the limit of factorial moment of $(L^{\prime}_{g,[a_{i},b_{i}]})_{i}$ is
$\lim_{g\to\infty}\mathbb{E}[(L^{\prime}_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L^{\prime}_{g,[a_{k},b_{k}]})_{n_{k}}]=\lim_{g\to\infty}\frac{\operatorname{Vol}({\mathcal{H}}_{g-K}(2^{g-1-K}))}{\operatorname{Vol}(\mathcal{C}^{\prime}_{g,Per}(2^{g-1}))}\prod^{k}_{i=1}[\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
From Proposition 5.1 and 5.2, we have
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}^{\prime}_{g,Per}(2^{g-1}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=1.$
From Theorem 2.2, we have
@@
$\lim_{g\to\infty}\mathbb{E}[(L^{\prime}_{g,[a_{1},b_{1}]})_{n_{1}}\cdots(L^{\prime}_{g,[a_{k},b_{k}]})_{n_{k}}]=\prod^{k}_{i=1}[3\pi(b_{i}^{2}-a_{i}^{2})]^{n_{i}}.$
By theorem 2.5, $(L^{\prime}_{g,[a_{1},b_{1}]},\cdots,L^{\prime}_{g,[a_{k},b_{k}]})$ converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a_{i},b_{i}]}$, where
$\lambda_{[a_{i},b_{i}]}=3\pi(b_{i}^{2}-a_{i}^{2}).$
From the property of Poisson distribution, we have
$\mathbb{P}(L^{\prime}_{g,[a_{1},b_{1}]}=n_{1},\cdots,L^{\prime}_{g,[a_{k},b_{k}]}=n_{k})=\prod_{i=1}^{k}\frac{\lambda_{[a_{i},b_{i}]}^{n_{i}}e^{-\lambda_{[a_{i},b_{i}]}}}{n_{i}!}.$
On the other hand
$\mathbb{P}(L^{\prime}_{g,[a_{1},b_{1}]}=n_{1},\cdots,L^{\prime}_{g,[a_{k},b_{k}]}=n_{k})=\frac{\operatorname{Vol}(\{(X,\omega):L^{\prime}_{g,[a_{i},b_{i}]}(X,\omega)=n_{i},i=1,\cdots,k\})}{\operatorname{Vol}(\mathcal{C}^{\prime}_{g,Per}(2^{g-1}))}.$
Note that
$\operatorname{Vol}(\{(X,\omega):L_{g,[a_{i},b_{i}]}(X,\omega)=n_{i},i=1,\cdots,k\})-\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))$
$\leq\operatorname{Vol}(\{(X,\omega):L^{\prime}_{g,[a_{i},b_{i}]}(X,\omega)=n_{i},i=1,\cdots,k\})$
$\leq\operatorname{Vol}(\{(X,\omega):L_{g,[a_{i},b_{i}]}(X,\omega)=n_{i},i=1,\cdots,k\}).$
From Proposition 5.1 and 5.2, we have
$\lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}^{\prime}_{g,Per}(2^{g-1}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=1,\\ \lim_{g\to\infty}\frac{\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}=0.$
So
$\mathbb{P}(L_{g,[a_{1},b_{1}]}=n_{1},\cdots,L_{g,[a_{k},b_{k}]}=n_{k})-\frac{\operatorname{Vol}(\mathcal{C}_{g}^{1}(\frac{B}{\sqrt{g}})\cup\mathcal{C}_{g}^{2}(\frac{B}{\sqrt{g}}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}$
$\leq\mathbb{P}(L^{\prime}_{g,[a_{1},b_{1}]}=n_{1},\cdots,L^{\prime}_{g,[a_{k},b_{k}]}=n_{k})\frac{\operatorname{Vol}(\mathcal{C}^{\prime}_{g,Per}(2^{g-1}))}{\operatorname{Vol}({\mathcal{H}}_{g}(2^{g-1}))}$
$\leq\mathbb{P}(L_{g,[a_{1},b_{1}]}=n_{1},\cdots,L_{g,[a_{k},b_{k}]}=n_{k}).$
Let $g\to\infty$ we have
$\mathbb{P}(L_{g,[a_{1},b_{1}]}=n_{1},\cdots,L_{g,[a_{k},b_{k}]}=n_{k})=\mathbb{P}(L^{\prime}_{g,[a_{1},b_{1}]}=n_{1},\cdots,L^{\prime}_{g,[a_{k},b_{k}]}=n_{k}).$
So $(L_{g,[a_{1},b_{1}]},\cdots,L_{g,[a_{k},b_{k}]})$ also converges jointly in distribution to a vector of random variables with Poisson distributions of means $\lambda_{[a_{i},b_{i}]}$, where
$\lambda_{[a_{i},b_{i}]}=3\pi(b_{i}^{2}-a_{i}^{2}).$
Since we don’t need to deal with simple zeros in the collapsing surgery, samilarly to the proof of theorem 1.2, we can prove Corollary 1.4.